# Rational Metareasoning for Large Language     Models

### v2

@@ Abstract
Being prompted to engage in reasoning has emerged as a core technique for using large language models (LLMs), deploying additional inference-time compute to improve task performance. However, as LLMs increase in both size and adoption, inference costs are correspondingly becoming increasingly burdensome. How, then, might we optimize reasoning’s cost-performance tradeoff? This work introduces a novel approach based on computational models of metareasoning used in cognitive science, training LLMs to selectively use intermediate reasoning steps only when necessary. We first develop a reward function that incorporates the Value of Computation by penalizing unnecessary reasoning, then use this reward function with Expert Iteration to train the LLM. Compared to few-shot chain-of-thought prompting and STaR, our method significantly reduces inference costs (20-37% fewer tokens generated across three models) while maintaining task performance across diverse datasets.
@@ 1 Introduction
Large language models (LLMs) rely on substantial computational power to handle complex problems (OpenAI et al., 2024; Chowdhery et al., 2022; de Vries, 2023). While initial studies mostly focused on the cost of training (Verdecchia et al., 2023), LLMs’ widespread deployment has made inference-time costs an increasingly important factor. Model compression techniques such as quantization, pruning, and knowledge distillation can lower post-training costs (Wan et al., 2024). However, there is a fundamental tension between inference cost and task performance: while many of these methods reduce costs at the expense of performance, others, such as chain-of-thought prompting (CoT; Wei et al., 2023; Kojima et al., 2023), do the opposite, raising inference costs to enhance task performance (Snell et al., 2024). It is worth noting that none of the previous approaches are adaptive: model compression modifications and existing CoT methods tend to raise or lower the inference cost on all queries, regardless of task complexity.
@@
In stark contrast to this static tradeoff, humans are able to adaptively allocate computational resources based on task difficulty (Krämer, 2014; Russell, 1997; Lieder & Griffiths, 2017). In this work, we draw inspiration from rational metareasoning – literally, reasoning about reasoning – a concept originally from the artificial intelligence literature (Russell & Wefald, 1991) that has been used to explain how humans adaptively manage computational resources (Lieder & Griffiths, 2017; Lieder et al., 2018; Griffiths et al., 2019a). Building on this, we develop a novel reward function based on the Value of Computation (VOC; Russell & Wefald, 1991), which formalizes the trade-off between inference cost and task performance. We adopt an iterative reinforcement learning process inspired by the Expert Iteration algorithm (Anthony et al., 2017). In each iteration, we generate multiple reasoning chains for each question. These reasoning chains are ranked using the reward function, and the dataset is filtered to retain only the best reasoning chain for each question. The model is then fine-tuned using this filtered dataset. However, unlike previous applications of Expert Iteration to LLMs (Zelikman et al., 2022), which filter generated examples based on correctness, our method optimizes for both the correctness and the cost of the reasoning process.
@@
We evaluated the effectiveness of our solution across a diverse set of tasks, from science knowledge (ARC; Clark et al., 2018) to commonsense reasoning (CommonsenseQA; Talmor et al., 2019), mathematical problem solving (GSM8K; Cobbe et al., 2021), and logical deductive reasoning (ProofWriter; Tafjord et al., 2021). Additionally, we assess the out-of-domain generalization on MMLU (Hendrycks et al., 2021), a multitask benchmark. Our approach achieves a substantial reduction in generated tokens (35.6-37.4% compared to few-shot prompting and 20.1-27.3% compared to STaR; Zelikman et al. 2022) while retaining comparable performance. Thus, we make the following contributions:
@@
    We introduce the idea of using rational metareasoning to optimize the tradeoff between inference cost and performance of LLMs.
@@
    We empirically demonstrate that rational metareasoning achieves the same task performance at lower inference costs (20-37% fewer tokens on average) across various datasets and reasoning tasks.
@@ 2 Rational Metareasoning
Unlike artificial intelligence, humans have limited time and cognitive resources (Griffiths et al., 2019b; Griffiths, 2020). We face diverse challenges requiring different approaches: avoiding a sudden obstacle when driving needs quick, intuitive thinking, whereas selecting a retirement investment strategy requires slow, deliberate reasoning (Krämer, 2014). Rational metareasoning (Russell & Wefald, 1991) suggests agents should adapt their reasoning based on the problem at hand. Intuitively, while reasoning solves a problem, metareasoning solves the problem of how to solve a problem: deciding which computations to perform while problem-solving. The essence of rational metareasoning is calculating the value of computation (VOC; Russell & Wefald, 1991) for each potential computation. The VOC balances the benefit of computation $c$ (characterized by the expected increase in the agent’s eventual utility) against its cost (usually time or energy).
To formalize this, agents are assumed to have some internal belief state $b\in\mathcal{B}$, which determines their expectation about the value of each action $a\in\mathcal{A}$: $\mathbb{E}[U(a)|b]$. A rational agent would simply choose the highest-value action: $a^{*}=\operatorname*{arg\,max}_{a\in\mathcal{A}}{[U(a)|b]}$. In contrast, a meta-rational agent can perform computation to change their belief state before choosing an action. Each computation $c\in\mathcal{C}$ updates the agent belief to $b^{\prime}$ with probability $P(b^{\prime}|c)$, which in turn affects their beliefs about the value of actions. However, each computation has an associated cost ($\text{cost}(c)$). The VOC then quantifies the value of performing computation $c$ given a starting belief state $b$,
@@
Thus, a meta-rational agent should pursue the computation $c^{*}$ with the highest VOC: $c^{*}=\operatorname*{arg\,max}_{c\in\mathcal{C}}VOC(c,b)$. If no computation has positive VOC, the agent should stop thinking and act in the world. Rational meta-reasoning can explain how humans allocate cognitive resources in various tasks (Lieder & Griffiths, 2017; Lieder et al., 2018; Callaway et al., 2018; 2021; 2022; Russek et al., 2022).
@@ 3 Rational Metareasoning with Large Language Models
To achieve an optimal balance between performance and efficiency, our approach introduces a new VOC-inspired reward function (Eq. 2) into an Expert Iteration training loop (Anthony et al., 2017; Zelikman et al., 2022), fine-tuning a LLM to produce reasoning chains adaptively depending on task difficulty.
@@ 3.1 Reward modeling
Chain-of-thought prompting is a technique in which a language model is encouraged to generate an intermediate output – a “chain of thought” – prior to producing the answer to a question (Wei et al., 2023; Kojima et al., 2023) We define the reward of a chain of thought as the difference between its utility and its cost,
@@
Specifically, $\pi_{\theta}(y|z,x)$ indicates the probability of generating the target sequence $y$ given both the chain of thought $z$ and the input $x$, while $\pi_{\theta}(y|x)$ denotes the probability of generating $y$ with only the input $x$. With respect to Equation 1, the language model’s initial belief about the value of actions or outputs is described by $\pi_{\theta}(y\mid x)$ (from Equation 3), whereas its final belief after generating a sequence of computations or tokens (a chain of thought $z$) is described by $\pi_{\theta}(y\mid z,x)$. The cost is directly proportional to the logarithm of the number of tokens in the chain of thought $l(z)$:
$\mathcal{C}(z)=\gamma\cdot\log l(z).$ (4)
The hyperparameter $\gamma$ scales the cost and utility to the same magnitude. We found that both linear and logarithmic cost functions perform adequately in our training algorithm, but the latter was better at balancing utility and cost, especially with higher sequence-length variance. A key benefit of this reward function is that it is parameterized by the same weights $\theta$ as the generative policy $\pi_{\theta}$, eliminating the need for an external reward model. This allows for direct estimation of the utility of a reasoning chain using the policy itself.
3.2 CoT Generation
We begin with a pretrained language model $\pi_{\theta}$ and an initial dataset of problems $x$ along with their corresponding correct final answers $y$: $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{D}$. Following prior work in online RL (Tang et al., 2024), we utilize the model itself to generate the reasoning chains, with few-shot prompting as a guide. Specifically, we concatenate a small set of examples, denoted as $\mathcal{P}$, each containing intermediate reasoning chains $z$, to each example in $\mathcal{D}$. For each task $\tau_{i}=(x_{i},y_{i})$ in the original dataset $\mathcal{D}$, we generate $K$ reasoning chains: $\hat{\tau}_{i}=\{(x_{i},z_{k,i},y_{i})\}_{k=1}^{K}$. If none of the $K$ generated reasoning chains for a task $\tau_{i}$ leads to the correct answer, we discard all samples. To minimize the likelihood of this outcome, we adopt the rationalization approach introduced by STaR (Zelikman et al., 2022): if the model fails to generate the correct answer, we generate a new reasoning chain by supplying the model with the correct answer in the prompt. Intuitively, rationalization increases the sample size, providing more options when selecting the $\operatorname*{arg\,max}$ of the VoC reward. We found that it provided marginal reduction in length with no effect on performance. Finally, we assess each reasoning chain using the Rational Metareasoning reward function (Equation 2).
@@ 3.3 Metareasoning Training
We demonstrate the effectiveness of our reward using a variation of the Expert Iteration algorithm (EI, Anthony et al., 2017). EI is known for its sample efficiency and strong performance on reasoning tasks (Havrilla et al., 2024; Zelikman et al., 2022). As an example of an online reinforcement learning (RL) algorithm, EI involves both exploration and policy improvement phases, with the policy $\pi_{\theta}$ being updated using data from the exploration phase. STaR (Zelikman et al., 2022) applies EI to LLMs by iteratively generating reasoning chains via few-shot prompting and compiling them into a fine-tuning dataset to refine the model. We build on this framework for our rational metareasoning training process, described in Algorithm 1.
@@
2:$\mathcal{D}_{0}\leftarrow\emptyset$
3:for $n$ in $1...N$ do $\triangleright$ Iterations
4:     ${\mathcal{D}_{n}}\leftarrow\{\mathcal{T}\subseteq\mathcal{D}\mid|\mathcal{T}|=T\}\cup\mathcal{D}_{n-1}$triangleright Sample batch from dataset
5:     for $k$ in $1...K$ do $\triangleright$ Perform reasoning chain generation
6:         $({z}_{i,k},{y}_{i,k})\leftarrow\pi_{n-1}(x_{i})\quad\forall i\in[1,D_{i}]$
7:         $({z}_{i,k},{y}_{i,k})\leftarrow\pi_{n-1}(\text{add\_hint}(x_{i}))\quad\forall i\in[1,D_{i}]\land y_{i,k}\neq y_{i}$triangleright Compute rationalization
8:     end for
9:     $r_{i,k}\leftarrow\mathcal{R}_{\pi_{n-1}}(x_{i},y_{i},z_{i,k})\quad\forall i,k(i\in[1,D_{i}]\land y_{i,k}=y_{i})$triangleright Compute reward
10:     $\hat{z}_{i}\leftarrow\arg\max_{k}\{r_{i,k}\}_{k=1}^{K}$triangleright Select best reasoning chain for each task $i$
11:     $\mathcal{D}_{n}^{*}\leftarrow\{(x_{i},\hat{z}_{i},y_{i})\in[1,D_{i}]\}$triangleright Create the optimal dataset
12:     $\pi_{n}\leftarrow\text{train}(\pi,\mathcal{D}_{n}^{*})$triangleright Finetune the original model on the optimal solutions
13:end for
Initially, in the exploration phase, we approximate the optimal policy $\hat{\pi}^{*}$ (whose computations maximize the VoC reward) by using rejection sampling on our student policy $\pi_{\theta}$. After generating $K$ intermediate reasoning chains $z_{1},\ldots,z_{K}$ for a given question $x$ (Section 3.2), we evaluate them using our reward function $\mathcal{R}_{\pi}$ (Section  3.1). We then construct $\mathcal{D}_{1}^{*}=\{(x_{i},z_{i},y_{i})\}_{i=1}^{N}$ by setting the rejection sampling threshold dynamically to the highest reward for each task $i$, therefore increasing robustness to hyperparameter choices : $z_{i}=\arg\max_{k}\{\mathcal{R}_{\pi}(x,z_{k},y)\}_{k=1}^{K}$. These rollouts are distilled into a policy $\pi_{1}$ using standard cross-entropy loss. This process can be iteratively repeated to refine the policy $\pi_{n}$ on the dataset $\mathcal{D}_{n}^{*}$. Finally, similar to STaR (Zelikman et al., 2024), instead of utilizing the entire training dataset at each iteration, as is typical in standard EI algorithms (Anthony et al., 2017; Gulcehre et al., 2023), we begin with a batch of $T$ steps during the first iteration and progressively increase the number of fine-tuning training steps by $T$ with each subsequent iteration. This approach allows the model to encounter new examples gradually, resulting in slower training initially, which ultimately enhances model performance.
@@ 4 Experiments
We now detail the datasets (Sec. 4.1), baselines (Sec. 4.2), and training (Sec 4.3) used to evaluate our method.
@@ 4.1 Datasets
In our efforts to develop a general-purpose reasoning model, we applied our training process and assessed its effectiveness across a diverse range of datasets and reasoning tasks. We constructed our training set by combining the training sets from these datasets into one dataset $\mathcal{D}$ and then evaluated the model on all corresponding test sets $\mathcal{T}$. Below is a detailed description of the datasets used:
  • ARC (Clark et al., 2018). The AI2 Reasoning Challenge (ARC) dataset comprises grade-school science questions, designed to evaluate a model’s capability to apply scientific knowledge.
  • CommonsenseQA (Talmor et al., 2019). This dataset is centered on commonsense question answering. It leverages implicit human knowledge that is commonly known and sensible, testing the model’s ability to provide answers based on everyday reasoning.
  • GSM8K (Cobbe et al., 2021). This dataset includes a variety of linguistically diverse grade school math word problems. It assesses the model’s proficiency in solving mathematical problems that require comprehension and application of arithmetic reasoning.
  • ProofWriter (Tafjord et al., 2021). This dataset assesses logical deductive reasoning by asking the model to determine if a conclusion can be inferred from premises presented in natural language.
These datasets have very different train split sizes. To ensure fairness and balance between the datasets, and to manage computational costs, we composed our training mixture by sampling 1,024 random samples from each of the training sets. We then evaluated the model on the public test set of each dataset. To further assess the generalization of our approach, we conducted out-of-distribution testing on MMLU (Hendrycks et al., 2021), a massive multitask benchmark consisting of multiple-choice questions from various branches of knowledge. Again, to limit computational costs, we conducted the test on the first 100 samples of each subject within the MMLU benchmark, giving us a total of 5700 questions across 57 subjects.

### v3

@@ Abstract
Recent approaches for leveraging large language models (LLMs) for reasoning often rely on additional inference-time computation to tackle complex tasks. However, as LLMs grow in size, the associated inference costs are becoming increasingly prohibitive. To address this, we introduce RaM (Rational Metareasoning): a cost-aware reasoning approach inspired by computational models of metareasoning in cognitive science. Our method trains LLMs to use intermediate reasoning steps selectively when they are likely to be beneficial. We first design a reward function that incorporates the Value of Computation by penalizing unnecessary reasoning, and then use it within an Expert Iteration framework to optimize the trade-off between performance and efficiency. Compared to few-shot chain-of-thought prompting and STaR, our method significantly reduces inference costs (23-45% fewer tokens generated across three models) while maintaining or increasing task performance across diverse datasets.
@@ 1 Introduction
Large language models (LLMs) rely on substantial computational power to handle complex problems (OpenAI et al., 2024; Chowdhery et al., 2022; de Vries, 2023). While initial studies mostly focused on the cost of training (Verdecchia et al., 2023), widespread use of LLMs has made inference-time costs an increasingly important factor. Moreover, there is a fundamental tension between inference cost and task performance: although chain-of-thought prompting (CoT; Wei et al., 2023; Kojima et al., 2023) and similar approaches improve task performance, they also substantially increase inference costs (Snell et al., 2024).
This trend has recently spiked with the development of Large Reasoning Models (OpenAI, 2024; DeepSeek-AI et al., 2025), instruction-tuned LLMs which are typically post-trained using reinforcement learning (RL) with outcome-based rewards to produce long chains of thought before each response. It is also worth noting that these approaches are not inherently adaptive: inference optimization (Wan et al., 2024) and existing CoT training methods often tend to raise or lower the inference cost on all queries, regardless of task complexity, or require the user to specify a budget up front (Anthropic et al., 2025).
In stark contrast to this static tradeoff, humans are able to adaptively allocate computational resources based on task difficulty (Kahneman, 2014; Russell, 1997; Lieder & Griffiths, 2017). In this work, we draw inspiration from rational metareasoning – literally, reasoning about reasoning – a concept originally from the artificial intelligence literature (Russell & Wefald, 1991) that has been used to explain how humans adaptively manage computational resources (Lieder & Griffiths, 2017; Lieder et al., 2018; Griffiths et al., 2019).
@@
Building on this, we develop a novel reward function based on the Value of Computation (VOC; Russell & Wefald, 1991), which formalizes the trade-off between inference cost and task performance. We adopt an iterative reinforcement learning process inspired by the Expert Iteration algorithm (Anthony et al., 2017). In each iteration, we generate multiple reasoning chains for each question. These reasoning chains are ranked using the reward function, and the dataset is filtered to retain only the best reasoning chain for each question. The model is then fine-tuned using this filtered dataset. However, unlike previous applications of Expert Iteration to LLMs (Zelikman et al., 2022; Havrilla et al., 2024), which filter generated examples solely based on the correctness of the final answer, our method optimizes for both correctness and the cost of the reasoning process.
@@
We evaluated the effectiveness of our solution across a diverse set of tasks, from scientific knowledge (ARC; Clark et al., 2018) to commonsense reasoning (CommonsenseQA; Talmor et al., 2019), mathematical problem solving (GSM8K; Cobbe et al., 2021), and logical deductive reasoning (ProofWriter; Tafjord et al., 2021). Additionally, we assess the out-of-domain generalization on MMLU (Hendrycks et al., 2021), a multitask benchmark. Our approach achieves a substantial reduction in generated tokens (35-42% compared to few-shot prompting, and 23-32% compared to STaR) while matching or increasing performance. Thus, we make the following contributions:
@@
    We employ rational metareasoning to optimize the tradeoff between inference cost and performance of LLMs on reasoning tasks.
@@
    We empirically demonstrate that rational metareasoning improves task performance at lower inference costs (23-42% fewer tokens on average) across various datasets and reasoning tasks.
@@ 2 Rational Metareasoning
Humans have limited time and cognitive resources (Griffiths et al., 2019; Griffiths, 2020). We face diverse challenges requiring different approaches: avoiding a sudden obstacle when driving needs quick, intuitive thinking, whereas selecting a retirement investment strategy requires slow, deliberate reasoning (Kahneman, 2014). Rational metareasoning (Russell & Wefald, 1991) captures this by suggesting agents should adapt their reasoning based on the problem at hand.
Intuitively, while reasoning solves a problem, metareasoning solves the problem of how to solve a problem: deciding which computations to perform while problem-solving. The essence of rational metareasoning is calculating the value of computation (VOC; Russell & Wefald, 1991) for each potential computation. The VOC balances the benefit of computation $c$ (the expected increase in the agent’s utility) against its cost (usually time or energy).
To formalize this, agents are assumed to have some internal belief state $b\in\mathcal{B}$, which determines their expectation about the value of each action $a\in\mathcal{A}$: $\mathbb{E}[U(a)|b]$. A rational agent would simply choose the highest-value action: $a^{*}=\text{argmax}_{a\in\mathcal{A}}{[U(a)|b]}$. However, this picture becomes more complex if an agent can perform computation to change their belief state before choosing an action. Each computation $c\in\mathcal{C}$ updates the agent’s belief to $b^{\prime}$ with probability $P(b^{\prime}|c)$, which in turn affects their beliefs about the value of actions, and incurs an associated cost ($\text{cost}(c)$). The VOC quantifies the value of performing computation $c$ given a starting belief state $b$,
@@
Thus, a meta-rational agent should pursue the computation $c^{*}$ with the highest VOC: $c^{*}=\text{argmax}_{c\in\mathcal{C}}VOC(c,b)$. If no computation has positive VOC, the agent should stop thinking and act in the world. Rational metareasoning can explain how humans allocate cognitive resources in various tasks (Lieder & Griffiths, 2017; Lieder et al., 2018; Callaway et al., 2018; 2021; 2022; Russek et al., 2022).
@@ 3 Rational Metareasoning with Large Language Models
To achieve an optimal balance between performance and efficiency, our approach introduces a new VOC-inspired reward function (Eq. 2) into Expert Iteration (Anthony et al., 2017; Zelikman et al., 2022), fine-tuning a LLM to produce reasoning chains adaptively.
Figure 1: RaM training. We iteratively generate reasoning chains using the current policy, score and filter them to approximate the optimal policy, and then finetune the base policy.
@@ 3.1 Reward modeling
Chain-of-thought prompting encourages LLMs to generate an intermediate output – a “chain of thought” – prior to producing the answer to a question (Wei et al., 2023; Kojima et al., 2023). We define the reward of a chain of thought as the difference between its utility and its cost,
@@
Specifically, $\pi_{\theta}(y|z,x)$ indicates the probability of generating the target sequence $y$ given both the chain of thought $z$ and the input $x$, while $\pi_{\theta}(y|x)$ denotes the probability of generating $y$ with only the input $x$. With respect to Equation 1, the language model’s initial belief about the value of actions or outputs is described by $\pi_{\theta}(y\mid x)$ (from Equation 3), whereas its final belief after generating a sequence of computations or tokens (a chain of thought $z$) is described by $\pi_{\theta}(y\mid z,x)$. The cost is directly proportional to the number of tokens in the chain of thought $l(z)$:
$\mathcal{C}(z)=\gamma\cdot l(z).$ (4)
The hyperparameter $\gamma$ scales the cost and utility to the same magnitude. A key benefit of this reward function is that it is parameterized by the same weights $\theta$ as the generative policy $\pi_{\theta}$, eliminating the need for an external reward model. This allows for direct estimation of the utility of a reasoning chain using the policy itself.
Finally, it’s worth noting that while standard reward models for RLHF bias the policy to produce longer reasoning chains (Singhal et al., 2024), thus necessitating appropriate corrections (Chen et al., 2024; Park et al., 2024) to correct for reward hacking, our reward model aims for and achieves the opposite effect.
@@ 3.2 Metareasoning Training
We demonstrate the effectiveness of our reward model using a variation of the Expert Iteration algorithm (EI, Anthony et al., 2017). EI is known for its sample efficiency and strong performance on reasoning tasks (Havrilla et al., 2024; Zelikman et al., 2022). As an example of an online reinforcement learning algorithm, EI involves both exploration and policy improvement phases, with the policy $\pi_{\theta}$ being updated using data from the exploration phase (as can be seen in Algorithm 1 and Figure 1).
Initially, in the exploration phase, we approximate the optimal policy $\hat{\pi}^{*}$ (whose computations maximize the VOC reward) by using rejection sampling on our student policy $\pi_{\theta}$. In particular, we begin with a pretrained language model $\pi_{\theta}$ and an initial dataset of problems $x$ along with their corresponding correct final answers $y$: $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{D}$. Following prior work in online RL (Tang et al., 2024), we utilize the model itself to generate the reasoning chains. Since we start with a pretrained model, we guide the model generation using few-shot prompting in the first iterations. Specifically, we concatenate a small set of examples, denoted as $\mathcal{P}$, each containing intermediate reasoning chains $z$, to each example in $\mathcal{D}$. During training and evaluation, we remove the few shot examples.
For each task $\tau_{i}=(x_{i},y_{i})$ in the original dataset $\mathcal{D}$, we generate $K$ reasoning chains: $\hat{\tau}_{i}=\{(x_{i},z_{k,i},y_{i})\}_{k=1}^{K}$. Then, we evaluate them using our reward function $\mathcal{R}_{\pi}$ (Section  3.1), and compute the advantage of each chain by subtracting the average reward for that question: $a_{i,k}=r_{i,k}-\frac{1}{K}\sum_{k^{\prime}=1}^{K}r_{i,k^{\prime}}$.
Using these advantage scores, we perform rejection sampling (by discarding reasoning chains with negative advantage) to construct the dataset of optimal trajectories $(\mathcal{D}_{1}^{*})$. Finally, we distill the selected rollouts into a policy $\pi_{1}$ via standard cross-entropy loss. This process can be iteratively repeated to refine the policy $\pi_{n}$ on the dataset $\mathcal{D}_{n}^{*}$. In closed-answer settings with verifiable answers, the reasoning chains that lead to incorrect answers can also be directly discarded, to produce a higher quality dataset.
@@
2:for $n$ in $1...N$ do
3:     for $k$ in $1...K$ do
4:         $({z}_{i,k},{y}_{i,k})\leftarrow\pi_{n-1}(x_{i})\quad\forall i\in[1,D]$triangleright Sample reasoning chains
5:     end for
6:     $r_{i,k}\leftarrow\mathcal{R}_{\pi_{n-1}}(x_{i},y_{i},z_{i,k})\quad\forall i,k(i\in[1,D],k\in[1,K])$triangleright Compute rewards
7:     $a_{i,k}\leftarrow r_{i,k}-\frac{1}{K}\sum_{k^{\prime}=1}^{K}r_{i,k^{\prime}}\quad\forall i,k(i\in[1,D],k\in[1,K])$triangleright Compute advantages
8:     $\hat{Z}_{i}\leftarrow\{\,z_{i,k}\mid a_{i,k}>0\,\}$triangleright Select best reasoning chains
9:     $\mathcal{D}_{n}^{*}\leftarrow\{\,(x_{i},\hat{z}_{i,k},y_{i})\mid\hat{z}_{i,k}\in\hat{Z}_{i},\,i\in[1,D]\,\}$triangleright Create the optimal dataset
10:     $\pi_{n}\leftarrow\text{train}(\pi,\mathcal{D}_{n}^{*})$triangleright Finetune the original model on the optimal solutions
11:end for
@@ 4.1 Datasets
To characterize the generality of our method, we applied it to a diverse range of datasets and reasoning tasks. We constructed our training set by combining the training sets from the following datasets into one dataset $\mathcal{D}$ and then evaluated the model on all corresponding public test sets $\mathcal{T}$. To facilitate comparison with baselines that use the correctness of the answer as a reward, we limited the training to datasets with verifiable answers, although this is not strictly required by our method. We used four datasets:
ARC (Clark et al., 2018). The AI2 Reasoning Challenge (ARC) dataset comprises grade-school science questions, designed to evaluate the capability to apply scientific knowledge.
CommonsenseQA (Talmor et al., 2019). This dataset is centered on commonsense question answering. It leverages implicit human knowledge and testing everyday reasoning.
GSM8K (Cobbe et al., 2021). This dataset includes a variety of linguistically diverse grade-school math word problems. It assesses proficiency in solving mathematical problems that require comprehension and application of arithmetic reasoning.
ProofWriter (Tafjord et al., 2021). This dataset assesses logical deductive reasoning by asking the model to determine if a conclusion follows from premises presented in natural language.
These datasets have very different train split sizes. To ensure fairness and balance between the datasets, and to manage computational costs, we composed our training mixture by sampling 1024 random samples from each of the training sets (for a total of 4096 samples). We then evaluated the model on the public test set of each dataset. To further assess the generalization of our approach, we conducted out-of-distribution testing on MMLU-CF (Zhao et al., 2024), a contamination free and more challenging variant of the original MMLU dataset (Hendrycks et al., 2021). This dataset consists of 10,000 multiple-choice questions from various branches of knowledge.